### v1

# LLM-Prompt: Integrated Heterogeneous Prompts for Unlocking LLMs in Time Series Forecasting

Abstract
Time series forecasting aims to model temporal dependencies among variables for future state inference, holding significant importance and widespread applications in real-world scenarios. Although deep learning-based methods have achieved remarkable progress, they still exhibit suboptimal performance in long-term forecasting and data-scarce scenarios. Recent research demonstrates that large language models (LLMs) achieve promising performance in time series forecasting. However, we find existing LLM-based methods still have shortcomings: (1) the absence of a unified paradigm for textual prompt formulation and (2) the neglect of modality discrepancies between textual prompts and time series. To address this, we propose LLM-Prompt, an LLM-based time series forecasting framework integrating multi-prompt information and cross-modal semantic alignment. Specifically, we first construct a unified textual prompt paradigm containing learnable soft prompts and textualized hard prompts. Second, to enhance LLMs’ comprehensive understanding of the forecasting task, we design a semantic space embedding and cross-modal alignment module to achieve cross-modal fusion of temporal and textual information. Finally, the transformed time series from the LLMs are projected to obtain the forecasts. Comprehensive evaluations on 6 public datasets and 3 carbon emission datasets demonstrate that LLM-Prompt is a powerful framework for time series forecasting.
1 Introduction
Time series forecasting, which utilizes past observations to predict future data, finds wide applications in domains such as electricity, weather, and carbon emissions [1]. Over the past five years, neural networks with diverse architectures—including linear models [2, 3], convolutional neural networks [4, 5], recurrent neural networks [6], and Transformers [7]—have been applied to time series forecasting. However, existing models face two critical challenges: On one hand, they rely heavily on large-scale labeled samples for parameter optimization, often suffering from limited generalization capability in cross-domain scenarios, which results in performance barriers under few-shot or zero-shot forecasting settings [8, 9]. On the other hand, existing methods struggle to capture long-term temporal dependencies, resulting in suboptimal long-term forecasting performance [10, 11].
Pretrained foundation models, such as Large Language Models (LLMs), have demonstrated strong few-shot and zero-shot learning capabilities in computer vision and natural language processing. With the technological advancements driven by LLMs, increasing research efforts focus on leveraging their potential for time series forecasting [12, 13, 14]. However, due to the modality gap between LLMs and time series data, knowledge transfer is required. As shown in Figure 1, the most common current approaches are: a) Time series-based LLMs: achieving modality alignment from time series to the semantic space of LLMs via learnable embedding layers [16]; b) Prompt-based LLMs: converting temporal data into textual modality through prompt engineering [9, 15]; c) Multimodal-based LLMs: using both time series and textual information as multimodal inputs, i.e. a hybrid architecture of the above two strategies [17, 18].
Figure 1: (a) Time series-based LLMs: Lack textual information, failing to leverage LLMs’ potential for time series forecasting. (b) Prompt-based LLMs: Struggle to capture fine-grained temporal patterns from textual prompts. (c) Multimodal-based LLMs: Use both text and time series as inputs but lack cross-modal alignment. (d) Our Method: Integrate multi-prompt information with time series through cross-modal alignment.
Current state-of-the-art approaches predominantly employ multimodal-based LLMs, yet notable limitations persist. On one hand, existing prompt engineering strategies utilize either hard prompts or soft prompts exclusively, failing to explore their synergy for further enhancing LLM performance [19, 20]. For hard prompts, they either solely utilize statistical information of time series or directly textualize raw time series [9]. Overall, a unified prompt paradigm remains absent. On the other hand, most works employ simple concatenation for multimodal data fusion, inadequately addressing the semantic gap and feature redundancy across different modalities [21, 22]. Notably, these technical limitations directly lead to significant shortcomings in long-term dependency modeling and noise robustness.
To address the aforementioned technical challenges, this paper proposes an LLM-empowered time series forecasting framework based on a collaborative mechanism of hard and soft prompts, termed LLM-Prompt. The core innovations of this framework lie in two aspects: a) constructing a unified paradigm for soft-hard prompt synergy and b) establishing a cross-modal information fusion network architecture. Specifically, we first design a dual-path prompt template integrating hard prompts (including temporal statistical features and raw numerical values) and adaptive soft prompts (optimized via gradient backpropagation) based on existing time series data. Then, we build a cross-modal data fusion module, including a semantic space embedding module and a cross-modal alignment module, to embed raw time series data and dual-path prompt texts into the LLMs’ dimension followed by modality adaptation. Finally, the LLMs’ outputs are projected to obtain time series forecasting results. Experimental results demonstrate that LLM-Prompt is effective for time series forecasting and verifies LLMs as capable few-shot and zero-shot time series learners. The main contributions of this study are summarized as follows:
  • We propose the LLM-Prompt framework, incorporating a soft-hard prompt synergy mechanism and cross-modal data fusion modules to address limitations in long-term dependency modeling and cross-domain generalization.
  • We develop a synergistic mechanism integrating hard prompts and soft prompts, advancing the establishment of a unified prompt engineering paradigm for time series.
  • We construct a semantic space embedding module and a cross-modal alignment module to enable cross-fusion of temporal data and textual information, enhancing feature representation.
  • Extensive experiments on 7 benchmark datasets and 3 carbon emission datasets demonstrate that LLM-Prompt achieves superior performance, particularly in few-shot and zero-shot scenarios.
2 Related Work
Deep learning has demonstrated superior performance in time series forecasting, primarily categorized into MLP-based, CNN-based, RNN-based, and Transformer-based architectures. Transformer-based models have emerged as the core research direction in recent years by capturing long-term dependencies and inter-variable correlations through self-attention mechanisms, achieving better predictive performance compared to other architectures. iTransformer treats multivariate time series as independent variable sequences, capturing cross-variable dependencies via attention mechanisms [23]. PatchTST divides time series into fixed-length patches, enhancing local pattern recognition through patch-level tokenization [24]. Furthermore, addressing the challenge of non-stationary time series, Non-stationary Transformer proposes series stationarization and de-stationary attention to mitigate non-stationarity in raw time series [25]. Other variants include Autoformer [26], which integrates decomposition frameworks with autocorrelation mechanisms, and Fedformer [27], which employs frequency-domain enhanced decomposition strategies, both significantly improving long-term forecasting accuracy.
Recent studies have challenged the effectiveness of Transformer-based models in time series forecasting. DLinear demonstrates that a simple linear regression in the original space achieves remarkable performance in both modeling and efficiency [2]. TimeMixer introduces a multi-scale mixing mechanism, enhancing forecasting through feature fusion at different temporal granularities [3]. CNN-based approaches, capturing local patterns via receptive fields, have also shown promising results in recent years. SCINet employs a hierarchical downsampling-convolution-interaction architecture to capture dynamic dependencies across temporal resolutions [28]. WaveNet’s dilated causal convolutions significantly expand the receptive field, addressing long-term dependency issues in traditional CNNs [29]. RNN-based models, limited by vanishing gradients in long-term forecasting tasks, have seen fewer research efforts recently. Nevertheless, Mamba—based on state space models—has reestablished an effective paradigm for time series forecasting [6]. This method dynamically selects critical temporal features to capture long-term dependencies. However, current deep learning frameworks remain constrained by sample size, exhibiting suboptimal performance in few-shot or zero-shot scenarios.
Recently, leveraging the cross-domain transfer capabilities of LLMs, they have demonstrated exceptional performance in time series forecasting tasks, significantly outperforming traditional methods, especially in few-shot scenarios [32]. As shown in Figure 1, the most prevalent current approaches include: a) Time series-based LLMs; b) Prompt-based LLMs; c) Multimodal-based LLMs. Time series-based LLMs achieve cross-modal alignment from time series to the LLM semantic space via learnable embedding layers, encoding temporal data into LLM-readable representations. Prompt-based LLMs convert numerical time series into textual information. Current research focuses on Multimodal-based LLMs, which perform cross-modal semantic alignment between temporal and textual data to overcome inherent limitations of single-modality representations. However, challenges remain in activating LLMs’ forecasting capabilities. Specifically, prompt engineering lacks a unified paradigm: TEMPO and S2IP-LLM solely employ learnable soft prompts, neglecting the importance of hard prompts [17, 20]; LLM-Time directly textualizes temporal data as hard prompts for LLMs, resulting in feature redundancy due to high overlap with raw temporal data [8]; Time-LLM is limited to constructing hard prompts using only statistical features of temporal data [19]. Furthermore, existing multimodal alignment methods rely on feature concatenation, failing to effectively bridge the modality gap [30, 31].
3 Methodology
3.1 Framework Overview
As illustrated in Figure 2, LLM-Prompt comprises three key modules: dual-path prompting, cross-modal data fusion (including semantic space embedding and cross-modal alignment), and time series forecasting.
Figure 2: LLM-Prompt framework. First, hard prompts and soft prompts are constructed based on time series. Second, the time series and prompts are embedded into the LLM’s semantic space, followed by cross-modal alignment. Finally, the fused information is fed into the pretrained LLM and projected through output layers to generate forecasting results.
3.2 Dual-Path Prompting
Soft Prompts. Soft prompts include prompt pool construction, similarity computation, and Top-k prompt selection. The core idea is to dynamically select soft prompt vectors most relevant to the time series [17, 20]. First, initialize a prompt pool. Then, compute the similarity between the time series data and prompt keys in the pool, selecting the Top-k most relevant prompts. Notably, the prompt pool is adaptively updated during iterations.
Prompt Pool Construction. Given prompt keys $K\in{R^{P\times D}}$ and prompt pool $V\in{R^{P\times L\times D}}$, where P is the pool size, L is the soft prompt length, and D is the embedding dimension. The prompt pool can be initialized through multiple strategies: zero initialization, uniform distribution initialization, or initialized from the vocabulary of pretrained LLMs. Notably, the prompt pool is updated during network iterations. Subsequent experiments further investigate the impact of pool size and soft prompt length on model performance.
Similarity Computation. Given the time series $X\in{R^{B\times N\times D}}$, where $B$ is the batch size, N is the variable dimension, and D is the embedding dimension. Compute the similarity matrix between X and K.
$S=Z{K^{T}}\in{R^{B\times N\times P}}$ (1)
Top-k Prompt Selection. Select the Top-k prompt key indices $I$ with the highest similarity. Retrieve the corresponding prompt vectors from the prompt pool using indices $I$ to construct soft prompts.
${\rm{softPromp}}{{\rm{t}}_{i}}=V[{I_{i}}]\in{R^{B\times N\times L\times D}}$ (2)
Figure 3: Hard prompt example
Hard Prompts. Hard prompts provide a direct yet effective approach to activate forecasting capabilities in LLMs. However, a well-considered unified paradigm remains absent. Based on our observation, existing works either solely textualize temporal data or construct text using only statistical features of time series, failing to integrate both approaches [19, 21]. Experimental practice indicates this limits performance gains. To address this, we propose a unified hard prompt paradigm. As illustrated in Figure 3, our paradigm comprises four key components: (1) Dataset Description, (2) Task Description, (3) Historical Data, and (4) Statistical Features. Dataset description provides background context for LLMs, while task description, historical data, and statistical features collectively enrich temporal information to enhance the LLMs’ forecasting capabilities. Subsequent experiments demonstrate that both auxiliary information [(1), (2), (3)] and temporal data [(4)] are all effective.
3.3 Cross-Modal Data Fusion
Semantic Space Embedding. This component primarily involves embedding time series data and hard prompts [21, 31, 22]. For soft prompts, they are generated directly in the embedding space and thus require no embedding.
Time Series. First, segment the time series into patches, then project them into the LLM’s embedding space to obtain $\hat{X}\in{R^{B\times N\times P\times D}}$, where P is the number of patches and $D$ is the embedding dimension. Given the pre-trained LLM’s vocabulary $E\in{R^{V\times D}}$, where $V$ is the vocabulary size, we realign time series data with the vocabulary through reprogramming, achieving textual representation of temporal data to enhance LLM comprehension. Specifically, learnable projection matrices generate query matrices $Q=\hat{X}{W^{Q}}$, key matrices $K=E{W^{K}}$, and value matrices $V=E{W^{V}}$ for each attention head. Then, multi-head self-attention produces the reprogramming temporal representations.
$\tilde{X}={\rm{MHSA}}(Q,K,V)\in{R^{B\times N\times P\times D}}$ (3)
Hard Prompt. Given the hard prompt $T$, map it to discrete tokens using the LLM’s pretrained tokenizer. Then, convert these discrete tokens into continuous vector representations via the LLM’s input embedding layer. During this process, both the tokenizer and LLM remain frozen with no parameter updates during training.
${\rm{hardPrompt}}={\rm{Embed}}({\rm{Tokenize}}(T))\in{R^{B\times N\times L\times D}}$ (4)
Cross-Modal Alignment. This module aims to achieve cross-modal alignment between soft-hard prompts and temporal data. Its core design resembles time series reprogramming in semantic space embedding. Specifically, cross-modal alignment is performed separately for soft/hard prompts and time series: (a) Dynamically align adaptive soft prompts with time series via a multi-head self-attention module, obtaining fused representation $Z_{1}$; (b) Process hard prompt text and time series using an identical architecture, obtaining fused representation $Z_{2}$. Finally, the joint representation $Z=[\tilde{X},Z_{1},Z_{2}]$ is formed through element-wise concatenation.
3.4 Time Series Forecasting
As shown in Figure 2, the joint representation $Z$ is fed into the frozen LLM. Subsequently, its output is passed to a projection layer to obtain the final prediction $Y$. GPT-2 is used by default unless otherwise specified, and subsequent experiments also investigate the impact of different LLM architectures on forecasting performance [33].
4 Experiments
To validate the effectiveness of LLM-Prompt, we conduct extensive experiments across various time series forecasting tasks, including long-term forecasting, few-shot forecasting, and zero-shot forecasting. Additionally, we investigate the positive effects of soft prompts and hard prompts in enhancing the forecasting capabilities of LLMs.
Baseline. We select 9 representative SOTAs from 5 categories: (1) LLMs-based: Time-LLM [19], TimeCMA [21]; (2) Transformer-based: Autoformer [26], Crossformer [35], iTransformer[23], PatchTST [36]; (3) CNN-based: TimeNet [5]; (4) RNN-based: Mamba[6]; (5) MLP-based: DLinear [2].
4.1 Long-term Forecasting
Setups. We conduct experiments on 9 datasets, including 6 classical benchmarks (ETTh1, ETTh2, ETTm1, ETTm2, exchange-rate, weather) and 3 carbon emission datasets (munichCarbon, chinaCarbon, worldCarbon). The input sequence length $T$ is fixed at 96, and we adopt four different forecasting horizons: $H\in\{60,90,120,150\}$ for chinaCarbon and worldCarbon, and $H\in\{96,192,336,720\}$ for the others. Evaluation metrics include Mean Squared Error (MSE) and Mean Absolute Error (MAE).
Results. The results summarized in Table 1, averaged over different forecasting horizons, demonstrate that LLM-Prompt outperforms all baselines in most cases. LLM-based methods surpass deep learning and linear models in long-term forecasting, aligning with one of this work’s motivations. Compared to other LLMs-based approaches, we observe that LLM-Prompt achieves a reduction of 3.5%/1.53% in average MSE/MAE over Time-LLM and 4.25%/3.07% over Time-CMA. This is attributed to our integration of both soft/hard prompts and cross-modal alignment in LLM-Prompt.
Table 1: Long-term forecasting results. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.433 0.438 $0.449$ 0.445 $0.449$ $0.447$ $0.520$ $0.492$ $0.460$ $0.455$ $0.460$ $0.457$ $0.477$ $0.475$ $0.549$ $0.530$ $0.464$ $0.455$ 0.446 $0.447$
ETTh2 0.369 0.398 0.380 0.403 $0.399$ $0.416$ $0.435$ $0.447$ $0.408$ $0.421$ $0.564$ $0.519$ $0.446$ $0.457$ $1.920$ $1.073$ $0.383$ $0.406$ $0.382$ $0.411$
ETTm1 0.381 0.399 $0.421$ $0.413$ $0.400$ $0.410$ $0.466$ $0.450$ $0.408$ $0.416$ $0.404$ $0.408$ $0.53$ $0.496$ $0.697$ $0.605$ $0.407$ $0.412$ 0.390 0.405
ETTm2 0.283 0.329 $0.291$ $0.335$ $0.294$ $0.336$ $0.331$ $0.365$ $0.299$ 0.333 $0.354$ $0.401$ $0.326$ $0.364$ $1.725$ $0.813$ $0.291$ $0.334$ 0.293 $0.334$
Exchange 0.355 0.400 $0.361$ 0.403 $0.448$ $0.451$ $0.761$ $0.584$ $0.425$ $0.447$ 0.340 $0.414$ $0.843$ $0.596$ $0.757$ $0.645$ $0.364$ $0.405$ $0.391$ $0.417$
Weather 0.254 0.279 $0.266$ $0.285$ $0.257$ $0.284$ $0.288$ $0.314$ $0.258$ $0.285$ $0.265$ $0.317$ $0.337$ $0.379$ $0.259$ $0.320$ $0.261$ $0.281$ 0.255 0.278
munichCarbon $0.412$ $0.465$ $0.411$ $0.465$ $0.433$ $0.486$ $0.485$ $0.518$ 0.415 0.469 $0.394$ $0.456$ $0.611$ $0.599$ 0.358 0.437 $0.412$ $0.465$ $0.430$ $0.477$
chinaCarbon 0.584 0.540 $0.627$ $0.557$ $0.659$ $0.586$ $0.660$ $0.592$ $0.665$ $0.588$ $0.689$ $0.642$ $0.642$ $0.608$ $1.177$ $0.834$ $0.667$ $0.579$ 0.604 0.546
worldCarbon $0.646$ $0.563$ $0.649$ $0.565$ 0.546 0.518 $0.687$ $0.614$ $0.751$ $0.639$ $0.754$ $0.698$ $0.653$ $0.603$ $0.619$ 0.539 $0.735$ $0.616$ 0.615 $0.547$
4.2 Few-shot Forecasting
Setups. LLMs demonstrate remarkable few-shot learning capabilities in CV and NLP. We further investigate whether LLMs exhibit similar capabilities in forecasting tasks. Most settings follow the long-term forecasting configuration. This section evaluates model performance under scenarios with only 10% and 5% of training samples.
Results. The results summarized in Tables 2 and 3, averaged over different forecasting horizons, demonstrate that LLMs-based methods significantly outperform deep learning models, demonstrating that LLMs retain few-shot learning capabilities in forecasting tasks. LLM-Prompt outperforms other LLMs-based approaches, which we attribute to the synergistic integration of soft/hard prompts better activating LLMs’ few-shot learning capabilities.
Table 2: Few-shot learning on 10% training data. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.411 0.452 0.477 $0.460$ $0.549$ $0.504$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ $0.510$ $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.477$ 0.455
ETTh2 0.383 0.404 $0.396$ $0.414$ $0.416$ $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ 0.391 0.412
ETTm1 0.411 0.408 $0.424$ $0.411$ $0.425$ $0.427$ $0.511$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ 0.414 0.406
ETTm2 0.288 0.332 $0.297$ $0.339$ $0.302$ $0.342$ $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ 0.290 0.333
Exchange 0.339 0.397 $0.368$ 0.407 $0.423$ $0.448$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.136$ $0.411$ $0.442$ $0.373$ $0.411$
Weather $0.270$ $0.289$ 0.269 0.288 0.255 0.282 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.339$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.273$ $0.289$
munichCarbon 0.412 0.465 $0.426$ $0.477$ $0.434$ $0.486$ $0.513$ $0.538$ $0.445$ $0.491$ $0.413$ $0.475$ $0.605$ $0.609$ 0.402 0.474 $0.443$ $0.491$ $0.423$ $0.474$
chinaCarbon 0.659 0.577 $0.714$ $0.606$ $0.690$ $0.596$ $0.785$ $0.651$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.043$ $0.833$ $0.792$ $0.645$ 0.665 0.580
worldCarbon $0.732$ $0.622$ $0.802$ $0.669$ 0.720 0.607 $0.950$ $0.733$ $0.959$ $0.740$ $0.845$ $0.699$ $0.781$ $0.659$ $0.966$ $0.782$ $0.937$ $0.727$ 0.725 0.609
Table 3: Few-shot learning on 5% training data. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.491 0.469 $0.527$ $0.491$ $0.583$ $0.521$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ 0.510 $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.518$ 0.479
ETTh2 0.388 0.408 $0.409$ $0.424$ 0.417 $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ $0.418$ 0.424
ETTm1 0.418 0.411 $0.425$ $0.413$ 0.435 0.431 $0.510$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ $0.433$ $0.419$
ETTm2 0.289 0.333 $0.294$ $0.337$ 0.304 0.343 $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ $0.303$ $0.348$
Exchange 0.366 0.410 $0.389$ $0.417$ $0.407$ $0.436$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.135$ $0.411$ $0.442$ $0.370$ 0.411
Weather 0.270 0.291 $0.273$ $0.293$ 0.264 0.290 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.338$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.280$ $0.296$
munichCarbon $0.419$ $0.472$ $0.437$ $0.487$ $0.435$ $0.486$ $0.513$ $0.537$ $0.445$ $0.491$ 0.402 0.469 $0.604$ $0.609$ 0.341 0.435 $0.443$ $0.491$ $0.426$ $0.476$
chinaCarbon 0.720 0.610 $0.743$ $0.623$ 0.721 0.612 $0.785$ $0.652$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.042$ $0.832$ $0.792$ $0.645$ $0.732$ $0.613$
worldCarbon 0.803 0.666 $0.859$ $0.703$ 0.745 0.631 $0.950$ $0.733$ $0.959$ $0.740$ $0.846$ $0.699$ $0.821$ $0.682$ $0.965$ $0.782$ $0.937$ $0.727$ $0.824$ $0.671$
4.3 Zero-shot Forecasting
Table 4: Zero-shot learning on carbon emission datasets. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE
$munich\rightarrow china$ 0.687 0.591 0.697 0.596 $0.731$ $0.607$ $0.797$ $0.633$
$munich\rightarrow world$ 0.897 0.705 0.934 0.719 $1.036$ $0.753$ $0.987$ $0.737$
$china\rightarrow munich$ 0.432 0.482 $0.449$ $0.495$ 0.446 $0.492$ $0.447$ 0.490
$china\rightarrow world$ 0.791 0.656 $0.818$ $0.664$ $0.876$ $0.694$ 0.740 0.613
$world\rightarrow munich$ 0.445 0.487 0.445 $0.492$ $0.448$ 0.491 $0.475$ $0.502$
$world\rightarrow china$ 0.631 $0.562$ $0.644$ 0.561 $0.689$ $0.587$ 0.616 0.545
Setups. LLMs possess the potential for zero-shot time series forecasting. We further investigate the cross-domain zero-shot learning capabilities of LLMs. Specifically, models are trained on one dataset and tested on another dataset with entirely unseen samples. Most experimental settings follow the long-term forecasting configuration. We evaluate zero-shot learning performance on ETT-series and carbon emission datasets. Due to differing variable dimensions in carbon emission datasets, only channel-independent models are evaluated.
Results. Brief results in Tables 4 and 5 show that LLMs-based methods and PatchTST outperform other deep learning approaches in cross-domain scenarios, demonstrating LLMs’ established zero-shot reasoning capability for time series forecasting. In the $ETTh1\rightarrow ETTm1$ setting, LLM-Prompt achieves reductions of 5.64% and 2.8% in MSE/MAE over Time-LLM and 10.1% and 5.64% over PatchTST.
Table 5: Zero-shot learning on ETT datasets. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM TimeCMA TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
$h1\rightarrow m1$ 96 0.723 0.553 $0.746$ 0.552 $0.772$ $0.565$ $0.766$ $0.573$ 0.713 $0.563$ $0.825$ $0.599$ $0.783$ $0.602$ $0.826$ $0.578$ $0.856$ $0.596$
192 $0.762$ 0.564 $0.832$ $0.587$ $0.807$ $0.581$ 0.749 0.568 0.737 $0.585$ $0.797$ $0.598$ $0.773$ $0.606$ $0.846$ $0.593$ $0.845$ $0.597$
336 0.755 0.557 $0.793$ 0.583 $0.807$ $0.586$ $0.843$ $0.589$ 0.757 $0.604$ $0.798$ $0.610$ $0.786$ $0.617$ $0.835$ $0.596$ $0.829$ $0.594$
720 0.765 0.584 $0.813$ 0.601 $0.819$ $0.601$ $0.865$ $0.605$ 0.800 $0.631$ $0.846$ $0.641$ $0.85$ $0.654$ $0.84$ $0.605$ $0.816$ $0.606$
Avg 0.751 0.564 $0.796$ 0.581 $0.801$ $0.583$ $0.806$ $0.584$ 0.752 $0.596$ $0.816$ $0.612$ $0.798$ $0.620$ $0.837$ $0.593$ $0.836$ $0.598$
$h2\rightarrow m2$ 96 0.213 0.301 0.221 $0.306$ $0.225$ $0.309$ $0.226$ $0.312$ $0.228$ $0.331$ $0.244$ $0.331$ $0.495$ $0.496$ $0.228$ $0.313$ 0.213 0.300
192 0.276 0.337 $0.279$ $0.337$ $0.286$ $0.341$ $0.287$ $0.343$ $0.300$ $0.379$ $0.298$ $0.358$ $0.638$ $0.562$ $0.287$ $0.344$ 0.273 0.334
336 0.330 0.365 0.335 0.370 $0.341$ $0.372$ $0.344$ $0.374$ $0.368$ $0.424$ $0.362$ $0.397$ $0.742$ $0.611$ $0.345$ $0.376$ $0.339$ $0.372$
720 0.432 0.420 $0.433$ 0.422 $0.438$ $0.426$ $0.443$ $0.431$ $0.514$ $0.507$ $0.470$ $0.456$ $1.017$ $0.741$ $0.442$ $0.427$ 0.429 0.420
Avg 0.313 0.356 $0.317$ $0.359$ $0.322$ $0.362$ $0.325$ $0.365$ $0.352$ $0.410$ $0.343$ $0.385$ $0.723$ $0.602$ $0.325$ $0.365$ 0.315 0.356
We further comprehensively investigate LLMs’ zero-shot learning capabilities for carbon emission forecasting through cross-domain experiments across three spatial scales (city, country, and global). Results demonstrate that LLM-Prompt frameworks serve as effective zero-shot predictors for carbon emissions, providing a technical foundation for achieving global carbon neutrality.
4.4 Ablation Study
We analyze the effectiveness of dual-path prompting and cross-modal alignment components. As shown in Table 7, ablations are conducted on multiple datasets (ETTm1, ETTh1, munichCarbon, chinaCarbon, worldCarbon). Removing either soft prompts, hard prompts, or cross-modal alignment degrades model performance. Prompt engineering is critical for large models to comprehend inputs: the absence of hard or soft prompts reduces performance by approximately 3.2%. The most severe impact occurs without cross-modal alignment, preventing the LLM from effectively understanding time series and causing a 3.7% performance drop.
Ablations on Hard Prompt. We further investigate how different types of information within hard prompts affect LLM-Prompt’s performance. As shown in Figure 3, we compare LLM-Prompt’s performance under three scenarios: (1) only auxiliary information, (2) only temporal data, and (3) full integration. As presented in Table 7, varying hard prompt types have minimal impact on the ETTh1 dataset but significantly affect the chinaCarbon dataset, where missing information causes 5̃performance degradation.
Table 6: Ablations on different components. Red: the best.
Methods LLM-Prompt w/o Soft Prompt w/o Hard Prompt w/o Cross-Modal Alignment
Metric MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.433 0.438 $0.444$ $0.442$ $0.443$ $0.441$ $0.443$ $0.441$
ETTm1 0.381 0.399 $0.402$ $0.408$ $0.403$ $0.409$ $0.410$ $0.409$
munichCarbon 0.412 0.465 $0.422$ $0.476$ $0.423$ $0.476$ $0.422$ $0.475$
chinaCarbon 0.584 0.540 $0.608$ $0.549$ $0.609$ $0.551$ $0.608$ $0.550$
worldCarbon 0.646 0.563 $0.660$ $0.572$ $0.658$ $0.575$ $0.665$ $0.577$
Table 7: Ablations on different hard prompt configurations. Red: the best.
Methods LLM-Prompt w/o Historical Data w/o Statistical Data
Metric MSE MAE MSE MAE MSE MAE
ETTh1 - 96 0.391 0.406 $0.389$ $0.406$ $0.389$ $0.406$
chinaCarbon - 96 0.472 0.477 $0.495$ $0.490$ $0.499$ $0.491$
Ablations on Soft Prompt. Additionally, we investigate the impact of soft prompt pool size and soft prompt length on LLM-Prompt’s performance. As shown in Figure 5, varying parameters have minimal effects on the ETTh1 dataset, while extreme parameter values (too large or too small) cause performance degradation on chinaCarbon. This may be attributed to information redundancy from oversized parameters or insufficient soft prompt information from undersized parameters.
Ablations on Parameter Sensitivity. We analyze the impact of intermediate projection dimension $d_{model}$ and multi-head attention head dimension $d_{ff}$ on model performance. As illustrated in Figure 5, varying parameter configurations have minimal impact on the ETTh1 dataset, possibly due to its sufficient sample size enabling effective model fitting. In contrast, for the chinaCarbon dataset, specific optimal configurations exist for both $d_{model}$ and $d_{ff}$, enabling optimal information flow within the network.
Figure 4: Ablations on different soft prompt parameters.
Figure 5: Ablations on different projection dimensions
5 Conclusion
This paper proposes LLM-Prompt to unlock the potential of LLMs in time series forecasting. The method incorporates a dual-path prompting mechanism, where the synergistic integration of dynamic selection of learnable soft prompts and semantic mapping of textualized hard prompts enhances LLMs’ comprehension of forecasting tasks. To further improve performance, a cross-modal data fusion architecture is designed through multi-head cross-attention modules, achieving alignment of textual and temporal features. Extensive experiments on multiple real-world datasets validate the effectiveness of LLM-Prompt in time series forecasting.
Despite initial progress, we observe that sole reliance on prompt engineering exhibits significant limitations in effectively activating LLMs’ forecasting capabilities—or, more critically, the intrinsic forecasting capability of LLMs remains unproven. We argue that this fundamental limitation stems from the text-centric training paradigm of LLMs, which inherently struggles to adapt to numerical tasks. Ours future work will explore integrating Chain-of-Thought techniques to unlock LLMs’ predictive potential or developing foundation models tailored for forecasting tasks.