### v3

# Time-Prompt: Integrated Heterogeneous Prompts for Unlocking LLMs in Time Series Forecasting

Abstract
Time series forecasting aims to model temporal dependencies among variables for future state inference, holding significant importance and widespread applications in real-world scenarios. Although deep learning-based methods have achieved remarkable progress, they still exhibit suboptimal performance in long-term forecasting. Recent research demonstrates that large language models (LLMs) achieve promising performance in time series forecasting, but this progress is still met with skepticism about whether LLMs are truly useful for this task. To address this, we propose Time-Prompt, a framework for activating LLMs for time series forecasting. Specifically, we first construct a unified prompt paradigm with learnable soft prompts to guide the LLM’s behavior and textualized hard prompts to enhance the time series representations. Second, to enhance LLM’ comprehensive understanding of the forecasting task, we design a semantic space embedding and cross-modal alignment module to achieve fusion of temporal and textual data. Finally, we efficiently fine-tune the LLM’s parameters using time series data. Furthermore, we focus on carbon emissions, aiming to provide a modest contribution to global carbon neutrality. Comprehensive evaluations on 6 public datasets and 3 carbon emission datasets demonstrate that Time-Prompt is a powerful framework for time series forecasting.
Introduction
Time series forecasting, which utilizes past observations to predict future data, finds wide applications in domains such as electricity, weather, and carbon emissions (Wang et al. 2024b). Over the past five years, neural networks with diverse architectures—including linear models (Zeng et al. 2023; Wang et al. 2024a), convolutional neural networks (Wang et al. 2023; Wu et al. 2023), recurrent neural networks (Gu and Dao 2023), and Transformers (Vaswani et al. 2017)—have been applied to time series forecasting. However, existing models face two critical challenges: On one hand, they rely heavily on large-scale labeled samples for parameter optimization, often suffering from limited generalization capability in cross-domain scenarios, which results in performance barriers under few-shot or zero-shot forecasting settings (Gruver et al. 2023; Liu et al. 2024a). On the other hand, existing methods struggle to capture long-term temporal dependencies, resulting in suboptimal long-term forecasting performance (Chen et al. 2024; Ni et al. 2025).
Pretrained foundation models, such as Large Language Models (LLMs), have demonstrated strong few-shot and zero-shot learning capabilities in computer vision and natural language processing. As pre-trained models advance, research is increasingly split between harnessing LLMs for time series forecasting (Shyalika et al. 2024; Jin et al. 2023; Zhang et al. 2024) and training time series foundation models (Goswami et al. 2024; Ansari et al. 2024; Wiliński et al. 2024). Considering the high computational cost of foundation models, our work focuses on the former. As shown in Figure 1, the most common LLMs-based approaches are: (a) Time series-based LLMs: Projecting the time series into the LLM’s semantic space via an embedding layer (Zhou et al. 2023); (b) Prompt-based LLMs: converting temporal data into textual modality through prompt engineering (Liu et al. 2024a); (c) Multimodal-based LLMs: Processing of time series and text separately (Zhou et al. 2025); (d) Multimodal-based LLMs: Feeding unified multimodal information into the LLMs (Zhao, Chen, and Sun 2025).
Figure 1: a) Time series-based LLMs: Lack textual information, failing to leverage LLMs’ potential for time series forecasting. (b) Prompt-based LLMs: Struggle to capture fine-grained temporal patterns from textual prompts. (c) Multimodal-based LLMs: Time series are processed via an encoder, and text via LLMs, separately. (d) Multimodal-based LLMs: Time series and text information are processed by LLMs in a unified manner.
Currently, research primarily focuses on two main types (c) and (d). For example, TimeCMA first encodes the multimodal data, where the time series is processed by an encoder and the text by a frozen LLM. Subsequently, cross-modal alignment is performed, and finally, the prediction result is obtained through a decoder (Liu et al. 2025a). In contrast, Time-LLM processes both time series and text uniformly through a frozen LLM, which then projects the target value (Jin et al. 2024).
Although current methods have achieved state-of-the-art performance, there remains some skepticism—do LLM truly benefit time series forecasting? We believe that with appropriate guidance, the time series forecasting potential of LLMs can be activated. This is because, for LLMs, textual reasoning is fundamentally a form of discrete-value reasoning. Viewed from this perspective, it is quite similar to the task of time series forecasting, with the main distinction being that time series consists of continuous values. Furthermore, LLMs have exhibited certain mathematical reasoning abilities, which also indicates a capacity for handling continuous values.
To unlock the potential of LLMs for time series forecasting, inspired by recent research (Liu et al. 2025a; Cao et al. 2024; Pan et al. 2024; Jin et al. 2024), we propose Time-Prompt, an LLM-empowered framework that activates the potential of LLMs for time series forecasting through time series reprogramming, dual-path prompting, cross-modal alignment, and LoRA. We design a dual-path prompting framework that activates the LLM’s specific patterns with soft prompts and enhances its time series representations with hard prompts. We also reprogram the time series to endow it with a text-like representation. Furthermore, we construct a cross-modal alignment module to eliminate the modality gap between text and time, and we efficiently fine-tune the LLM with LoRA to enhance its ability to process continuous values. In addition, given that carbon emissions have become a globally critical issue, we conduct a cross-scale study on carbon emission forecasting—spanning global, national, and urban levels—to provide technical support for the strategic requirements of global carbon neutrality goals. The main contributions of this study are summarized as follows:
  • We propose Time-Prompt, a framework that activates the potential of LLMs in time series forecasting through strategies such as dual-path prompting, cross-modal alignment, and efficient fine-tuning.
  • We design a dual-path prompting strategy, with learnable soft prompts guiding the LLM’s behavior and hard prompts extracting enhanced time series representations.
  • We construct a semantic space embedding module and a cross-modal alignment module to enable cross-modal fusion of temporal data and textual information, eliminating the modality gap.
  • Extensive experiments on 6 benchmark datasets and 3 carbon emission datasets demonstrate that Time-Prompt achieves superior performance.
Related Work
Deep learning has demonstrated superior performance in time series forecasting, primarily categorized into MLP-based, CNN-based, RNN-based, and Transformer-based architectures. Transformer-based models have emerged as the core research direction in recent years by capturing long-term dependencies and inter-variable correlations through self-attention mechanisms, achieving better predictive performance compared to other architectures. iTransformer treats multivariate time series as independent variable sequences, capturing cross-variable dependencies via attention mechanisms (Liu et al. 2023). PatchTST divides time series into fixed-length patches, enhancing local pattern recognition through patch-level tokenization (Nie et al. 2023). Furthermore, addressing the challenge of non-stationary time series, Non-stationary Transformer proposes series stationarization and de-stationary attention to mitigate non-stationarity in raw time series (Liu et al. 2022b). Other variants include Autoformer (Wu et al. 2021), which integrates decomposition frameworks with autocorrelation mechanisms, and Fedformer (Zhou et al. 2022), which employs frequency-domain enhanced decomposition strategies, both significantly improving long-term forecasting accuracy.
Figure 2: Time-Prompt framework. First, hard prompts and soft prompts are constructed based on time series. Second, the time series and prompts are embedded into the LLM’s semantic space, followed by cross-modal alignment. Finally, the fused information is fed into the pretrained LLM and projected through output layers to generate forecasting results.
Recent studies have challenged the effectiveness of Transformer-based models in time series forecasting. DLinear demonstrates that a simple linear regression in the original space achieves remarkable performance in both modeling and efficiency (Zeng et al. 2023). TimeMixer introduces a multi-scale mixing mechanism, enhancing forecasting through feature fusion at different temporal granularities (Wang et al. 2024a). CNN-based approaches, capturing local patterns via receptive fields, have also shown promising results in recent years. SCINet employs a hierarchical downsampling-convolution-interaction architecture to capture dynamic dependencies across temporal resolutions (Liu et al. 2022a). WaveNet’s dilated causal convolutions significantly expand the receptive field, addressing long-term dependency issues in traditional CNNs (Van Den Oord et al. 2016). RNN-based models, limited by vanishing gradients in long-term forecasting tasks, have seen fewer research efforts recently. Nevertheless, Mamba—based on state space models—has reestablished an effective paradigm for time series forecasting (Gu and Dao 2023). This method dynamically selects critical temporal features to capture long-term dependencies. However, current deep learning frameworks remain constrained by sample size, exhibiting suboptimal performance in few-shot or zero-shot scenarios.
Recently, leveraging the cross-domain transfer capabilities of LLMs, they have demonstrated exceptional performance in time series forecasting tasks, significantly outperforming traditional methods, especially in few-shot scenarios (Liu et al. 2024b). As shown in Figure 1, the most prevalent current approaches include: (a) Time series-based LLMs; (b) Prompt-based LLMs; (c) Multimodal-based LLMs; (d) Multimodal-based LLMs. Time series–based LLMs project time series into the semantic space of LLMs via an embedding layer, and subsequently map the embedding vectors to obtain prediction values (Zhou et al. 2023). In contrast, Prompt-based LLMs construct the time series as a text modality (Xue and Salim 2023). Lstprompt is a typical Prompt-based method that proposes a prompt construction technique, guiding the LLM to make predictions via a Chain-of-Thought approach (Liu et al. 2024a).
However, single-modality time series forecasting often suffers from limited representational capacity, leading to suboptimal prediction performance. Therefore, many studies attempt to integrate Type (a) and Type (b) approaches, leveraging multimodal information to overcome the inherent limitations of single-modality methods. Early efforts predominantly fall into Type (d), which uniformly feeds time series and textual inputs into LLMs to obtain improved feature representations. GPT4MTS, a foundational work in this category, embeds time series and simple text prompts into the semantic space, from which the large model derives a joint representation (Jia et al. 2024). Many subsequent works build upon this framework. Time-LLM enhances model comprehension of time series by reprogramming the input time series (Jin et al. 2024). Meanwhile, S²IP-LLM and TEMPO employ learnable soft prompts to mark specific temporal patterns, thereby more effectively activating the large model (Cao et al. 2024; Pan et al. 2024).
Subsequently, with growing skepticism about whether LLMs are truly useful for time series (Tan et al. 2024), some studies shift away from feeding time series into LLMs for representation and instead adopt dedicated encoders—this approach corresponds to Type (c). TimeCMA encodes the time series with an encoder and the text with an LLM, followed by cross-modal alignment, and finally generates the prediction via a unified decoder (Liu et al. 2025a). In a related approach, another study further enhances the textual information using reinforcement learning (Su et al. 2025).
The problem with Type (c) is that the LLM only plays an auxiliary role, contributing little to time series forecasting. The issue with Type (d), however, is the LLM’s inability to handle continuous values like those in time series. Nevertheless, recent research continues to emerge for both Type (c) (Zhou et al. 2025; Su et al. 2025) and Type (d) (Chen et al. 2025; Zhang et al. 2025), in an effort to address the widespread skepticism. Our method belongs to the Type (d). LLMs possess powerful text-based reasoning capabilities. Fundamentally, their primary distinction from time series lies in the nature of data: discrete versus continuous values. Therefore, time series forecasting and textual reasoning exhibit significant similarity. Moreover, LLMs have also demonstrated strong reasoning abilities in mathematics. So we believe that with appropriate guidance, LLMs can enhance the performance of time series forecasting. Our method mainly contributes in this area. On the LLM side, we guide the LLM’s latent patterns with soft prompts and achieve efficient fine-tuning through LoRA. For the time series aspect, we reprogram the time series data and incorporate cross-modal alignment.
Methodology
Framework Overview
The overall framework of Time-Prompt is illustrated in Figure 2. First, we construct a dual-path prompting framework for the time series forecasting task. This framework incorporates both hard prompts (including temporal statistical features and historical data) and learnable soft prompts. Our goal is to guide the LLM’s internal behavioral patterns via soft prompts and enhance time series feature representation through hard prompts. Second, we embed both the time series and text into the LLM’s semantic space via semantic space embedding module to facilitate better comprehension by the LLM. Subsequently, to address the modality gap between text and time series, we employ cross-modal alignment. Finally, the global and local features of the time series are extracted and projected to obtain the predicted values. Here, the LLM is fine-tuned using LoRA to adapt it from discrete-value reasoning to continuous-value prediction.
Dual-Path Prompting
Soft Prompts. Soft prompts include prompt pool construction, similarity computation, and Top-k prompt selection. The core idea is to dynamically select soft prompt vectors most relevant to the time series. First, initialize a prompt pool. Then, compute the similarity between the time series data and prompt keys in the pool, selecting the Top-k most relevant prompts. Notably, the prompt pool is adaptively updated during iterations.
Prompt Pool Construction. Given prompt keys $K\in{R^{P\times D}}$ and prompt pool $V\in{R^{P\times L\times D}}$, where P is the pool size, L is the soft prompt length, and D is the embedding dimension. The prompt pool can be initialized through multiple strategies: zero initialization, uniform distribution initialization, or initialized from the vocabulary of pretrained LLMs. Notably, the prompt pool is updated during backpropagation. Subsequent experiments further investigate the impact of pool size and soft prompt length on model performance.
Similarity Computation. Given the time series $X\in{R^{B\times N\times D}}$, where $B$ is the batch size, $N$ is the variable dimension, and $D$ is the embedding dimension. Compute the similarity matrix between X and K.
$S=Z{K^{T}}\in{R^{B\times N\times P}}$ (1)
Top-k Prompt Selection. Select the Top-k prompt key indices $I$ with the highest similarity. Retrieve the corresponding prompt vectors from the prompt pool using indices $I$ to construct soft prompts.
${\rm{softPromp}}{{\rm{t}}_{i}}=V[{I_{i}}]\in{R^{B\times N\times L\times D}}$ (2)
Figure 3: Hard prompt example on chinaCarbon dataset.
Hard Prompts. Hard prompts provide a direct and effective way to activate the LLM’s time series forecasting capability, as no modality gap exists between the LLM and the prompts. Hard prompts can also be viewed as enhanced representations of the time series. As illustrated in Figure 3, our paradigm comprises three key components: (1) Instruction, (2) Statistical Characteristics, and (3) Historical Data. The Instruction provides the LLM with the task context, the Statistical characteristics correctly guide the LLM’s patterns, and the Historical Data is the direct textualization of the time series data.
Table 1: Long-term forecasting results. Bold: the best, Underline: the second best.
Methods TSF-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.430 0.439 $0.449$ 0.445 $0.449$ $0.447$ $0.520$ $0.492$ $0.460$ $0.455$ $0.460$ $0.457$ $0.477$ $0.475$ $0.549$ $0.530$ $0.464$ $0.455$ 0.446 $0.447$
ETTh2 0.368 0.397 0.380 0.403 $0.399$ $0.416$ $0.435$ $0.447$ $0.408$ $0.421$ $0.564$ $0.519$ $0.446$ $0.457$ $1.920$ $1.073$ $0.383$ $0.406$ $0.382$ $0.411$
ETTm1 0.380 0.398 $0.421$ $0.413$ $0.400$ $0.410$ $0.466$ $0.450$ $0.408$ $0.416$ $0.404$ $0.408$ $0.53$ $0.496$ $0.697$ $0.605$ $0.407$ $0.412$ 0.390 0.405
ETTm2 0.281 0.326 $0.291$ $0.335$ $0.294$ $0.336$ $0.331$ $0.365$ $0.299$ 0.333 $0.354$ $0.401$ $0.326$ $0.364$ $1.725$ $0.813$ $0.291$ $0.334$ 0.293 $0.334$
Exchange 0.331 0.390 $0.361$ 0.403 $0.448$ $0.451$ $0.761$ $0.584$ $0.425$ $0.447$ 0.340 $0.414$ $0.843$ $0.596$ $0.757$ $0.645$ $0.364$ $0.405$ $0.391$ $0.417$
Weather 0.252 0.277 $0.266$ $0.285$ $0.257$ $0.284$ $0.288$ $0.314$ $0.258$ $0.285$ $0.265$ $0.317$ $0.337$ $0.379$ $0.259$ $0.320$ $0.261$ $0.281$ 0.255 0.278
munichCarbon 0.411 0.464 $0.411$ $0.465$ $0.433$ $0.486$ $0.485$ $0.518$ $0.415$ $0.469$ $0.394$ $0.456$ $0.611$ $0.599$ 0.358 0.437 $0.412$ $0.465$ $0.430$ $0.477$
chinaCarbon 0.571 0.532 $0.627$ $0.557$ $0.659$ $0.586$ $0.660$ $0.592$ $0.665$ $0.588$ $0.689$ $0.642$ $0.642$ $0.608$ $1.177$ $0.834$ $0.667$ $0.579$ 0.604 0.546
worldCarbon $0.617$ $0.544$ $0.649$ $0.565$ 0.546 0.518 $0.687$ $0.614$ $0.751$ $0.639$ $0.754$ $0.698$ $0.653$ $0.603$ $0.619$ 0.539 $0.735$ $0.616$ 0.615 $0.547$
Table 2: Few-shot learning on 10% training data. Bold: the best, Underline: the second best.
Methods TSF-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.452 0.445 0.477 $0.460$ $0.549$ $0.504$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ $0.510$ $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.477$ 0.455
ETTh2 0.381 0.404 $0.396$ $0.414$ $0.416$ $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ 0.391 0.412
ETTm1 0.394 0.402 $0.424$ $0.411$ $0.425$ $0.427$ $0.511$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ 0.414 0.406
ETTm2 0.284 0.327 $0.297$ $0.339$ $0.302$ $0.342$ $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ 0.290 0.333
Exchange 0.337 0.395 $0.368$ 0.407 $0.423$ $0.448$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.136$ $0.411$ $0.442$ $0.373$ $0.411$
Weather 0.265 0.286 $0.269$ $0.288$ 0.255 0.282 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.339$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.273$ $0.289$
munichCarbon 0.406 0.459 $0.426$ $0.477$ $0.434$ $0.486$ $0.513$ $0.538$ $0.445$ $0.491$ $0.413$ $0.475$ $0.605$ $0.609$ 0.402 0.474 $0.443$ $0.491$ $0.423$ $0.474$
chinaCarbon 0.653 0.575 $0.714$ $0.606$ $0.690$ $0.596$ $0.785$ $0.651$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.043$ $0.833$ $0.792$ $0.645$ 0.665 0.580
worldCarbon 0.705 0.601 $0.802$ $0.669$ 0.720 0.607 $0.950$ $0.733$ $0.959$ $0.740$ $0.845$ $0.699$ $0.781$ $0.659$ $0.966$ $0.782$ $0.937$ $0.727$ $0.725$ $0.609$
Cross-Modal Data Fusion
Semantic Space Embedding. This component primarily involves embedding the time series and prompts into the LLM’s semantic space to enhance the LLM’s understanding of the variables. For soft prompts, they are generated directly in the embedding space and thus require no embedding.
Time Series. First, segment the time series into patches, then project them into the LLM’s embedding space to obtain $\hat{X}\in{R^{B\times N\times P\times D}}$, where P is the number of patches and $D$ is the embedding dimension. Given the pre-trained LLM’s vocabulary $E\in{R^{V\times D}}$, where $V$ is the vocabulary size, we realign time series data with the vocabulary through reprogramming, achieving textual representation of temporal data to enhance LLM comprehension. Specifically, learnable projection matrices generate query matrices $Q=\hat{X}{W^{Q}}$, key matrices $K=E{W^{K}}$, and value matrices $V=E{W^{V}}$ for each attention head. Then, multi-head cross-attention produces the reprogramming temporal representations.
$\tilde{X}={\rm{MHCA}}(Q,K,V)\in{R^{B\times N\times P\times D}}$ (3)
Hard Prompt. Given the hard prompt $T$, map it to discrete tokens using the LLM’s pretrained tokenizer. Then, convert these discrete tokens into continuous vector representations via the LLM’s input embedding layer. During this process, the LLM’s tokenizer remain frozen with no parameter updates during training.
${\rm{hardPrompt}}={\rm{Embed}}({\rm{Tokenize}}(T))\in{R^{B\times N\times L\times D}}$ (4)
Cross-Modal Alignment. This module is designed to achieve cross-modal fusion among the time series data, soft prompts, and hard prompts to eliminate the modality gap. Many existing works simply concatenate these variables, which can lead to severe feature redundancy when the number of prompt tokens is large (Jin et al. 2024). Our specific approach is as follows: (a) We use a multi-head cross-attention mechanism to align the soft prompt with the time series, obtaining a fused representation $Z_{1}$. Similarly, the hard prompt is aligned with the time series to obtain $Z_{2}$. (b) We then fuse the original time series and the fused representations $Z_{1}$ and $Z_{2}$ via a gating mechanism, yielding the final representation $Z$.
$Z=g_{0}\odot\tilde{X}+g_{1}\odot Z_{1}+g_{2}\odot Z_{2}$ (5)
Table 3: Few-shot learning on 5% training data. Bold: the best, Underline: the second best.
Methods TSF-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.464 0.454 $0.527$ $0.491$ $0.583$ $0.521$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ 0.510 $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.518$ 0.479
ETTh2 0.387 0.406 $0.409$ $0.424$ 0.417 $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ $0.418$ 0.424
ETTm1 0.400 0.404 $0.425$ $0.413$ 0.435 0.431 $0.510$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ $0.433$ $0.419$
ETTm2 0.285 0.329 $0.294$ $0.337$ 0.304 0.343 $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ $0.303$ $0.348$
Exchange 0.358 0.403 $0.389$ $0.417$ $0.407$ $0.436$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.135$ $0.411$ $0.442$ $0.370$ 0.411
Weather 0.268 0.288 $0.273$ $0.293$ 0.264 0.290 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.338$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.280$ $0.296$
munichCarbon $0.416$ $0.469$ $0.437$ $0.487$ $0.435$ $0.486$ $0.513$ $0.537$ $0.445$ $0.491$ 0.402 0.469 $0.604$ $0.609$ 0.341 0.435 $0.443$ $0.491$ $0.426$ $0.476$
chinaCarbon 0.654 0.577 $0.743$ $0.623$ 0.721 0.612 $0.785$ $0.652$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.042$ $0.832$ $0.792$ $0.645$ $0.732$ $0.613$
worldCarbon 0.738 0.619 $0.859$ $0.703$ 0.745 0.631 $0.950$ $0.733$ $0.959$ $0.740$ $0.846$ $0.699$ $0.821$ $0.682$ $0.965$ $0.782$ $0.937$ $0.727$ $0.824$ $0.671$
Table 4: Ablation Experiment
Methods Time-Prompt w/o SP w/o HP w/o CMA w/o LoRA Qwen 3 0.6B
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.430 0.439 $0.434$ $0.441$ $0.437$ $0.442$ $0.434$ $0.442$ $0.435$ $0.440$ $0.717$ $0.579$
ETTm1 0.380 0.398 $0.383$ $0.399$ $0.384$ $0.400$ $0.386$ $0.402$ $0.386$ $0.401$ $0.719$ $0.564$
munichCarbon 0.411 0.532 $0.413$ $0.466$ $0.414$ $0.467$ $0.414$ $0.467$ $0.416$ $0.468$ $0.652$ $0.588$
chinaCarbon 0.571 0.532 $0.581$ $0.540$ $0.584$ $0.540$ $0.587$ $0.544$ $0.583$ $0.538$ $0.802$ $0.652$
worldCarbon 0.617 0.544 $0.636$ $0.558$ $0.633$ $0.556$ $0.633$ $0.556$ $0.630$ $0.552$ $0.869$ $0.713$
Time Series Forecasting
Many existing works rely solely on frozen LLMs. Given the high dimensionality of the LLM’s semantic space, they often resort to direct truncation to prevent the curse of dimensionality. This approach, however, is detrimental to activating the LLM’s potential for time series forecasting. In this module, we first employ LoRA to efficiently fine-tune the LLM, enhancing its ability to comprehend continuous values. Second, to address the excessively high semantic dimensionality, we project its representations into a lower-dimensional space, extract global features, and then jointly project these global features together with local features to produce the final prediction $Y$.
$F_{\text{local}}$ $=\text{Linear}(Z),$ (6)
$F_{\text{global}}$ $=\text{Mean}(F_{\text{local}}),$ (7)
$F_{\text{fused}}$ $=\text{Linear}([F_{\text{local}};F_{\text{global}}]),$ (8)
$Y$ $=\text{Head}(\text{Mean}(F_{\text{fused}})).$ (9)
Experiments
Dataset. We utilize six widely-used public datasets covering multiple domains, including electricity, finance, and weather, as well as three carbon emission datasets with different spatial scales. Among them, the Munich carbon emission dataset was collected by one of our co-authors. We conduct extensive experiments across these nine datasets to demonstrate the superiority of Time-Prompt, including tasks such as long-term forecasting and few-shot forecasting. Zero-shot forecasting is not considered in this work, as we think LLM-based frameworks currently lack strong generalization capability in such settings.
Baseline. We select 9 representative SOTAs from 5 categories: (1) LLMs-based: Time-LLM (Jin et al. 2024), TimeCMA (Liu et al. 2025a); (2) Transformer-based: Autoformer (Wu et al. 2021), Crossformer (Zhang and Yan 2023), iTransformer(Liu et al. 2023), PatchTST (Nie et al. 2022); (3) CNN-based: TimeNet (Wu et al. 2023); (4) RNN-based: Mamba(Gu and Dao 2023); (5) MLP-based: DLinear (Zeng et al. 2023).
Implementation. Since the vast majority of LLM-based time series forecasting methods adopt GPT-2, we also use GPT-2 by default in this work for fair comparison (Radford et al. 2019). The evaluation metrics are Mean Squared Error (MSE) and Mean Absolute Error (MAE). Additional details (e.g., datasets, baseline, full results) are provided in the appendix.
Long-term Forecasting
Setups. We conduct experiments on 9 datasets, including 6 classical benchmarks (ETTh1, ETTh2, ETTm1, ETTm2, exchange-rate, weather) and 3 carbon emission datasets (munichCarbon, chinaCarbon, worldCarbon). The input sequence length $T$ is fixed at 96, and we adopt four different forecasting horizons: $H\in\{60,90,120,150\}$ for chinaCarbon and worldCarbon, and $H\in\{96,192,336,720\}$ for the others.
Results. The results summarized in Table 1, averaged over different forecasting horizons, demonstrate that Time-Prompt outperforms all baselines in most cases. Time-Prompt achieves the best performance in 14 out of 18 benchmark settings. Overall, LLM-based methods demonstrate superior performance compared to conventional deep learning approaches. Specifically, Time-Prompt reduces MSE and MAE by 5.5% and 2.7% over Time-LLM, by 6.3% and 4.2% over TimeCMA, and by 4.3% and 2.4% compared to PatchTST—the strongest deep learning baseline.
Few-shot Forecasting
Setups. LLMs demonstrate remarkable few-shot learning capabilities in CV and NLP. We further investigate whether LLMs exhibit similar capabilities in forecasting tasks. Most settings follow the long-term forecasting configuration. This section evaluates model performance under scenarios with only 10% and 5% of training samples.
Results. The brief results summarized in Tables 2 and 3, averaged over different forecasting horizons, demonstrate that LLMs-based methods significantly outperform deep learning models, demonstrating that LLMs retain few-shot learning capabilities in forecasting tasks. The superiority of Time-Prompt is even more pronounced in few-shot scenarios compared to long-term forecasting. Time-Prompt achieves the best performance in 30 out of 36 benchmark settings Under the 10% training data setting, Time-Prompt reduces MSE and MAE by 7.1% and 4.3% over Time-LLM, and by 8.0% and 5.5% over TimeCMA. Under the more challenging 5% training data setting, it further achieves 8.8% and 5.7% MSE/MAE reductions compared to Time-LLM, and 7.9% and 5.5% reductions compared to TimeCMA.
Table 5: Modal Analysis.
soft prompt 3 5 10 15
MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.376 0.402 0.376 0.402 0.374 0.401 0.377 0.402
munichCarbon 0.238 0.352 0.239 0.352 0.237 0.351 0.236 0.351
chinaCarbon 0.473 0.472 0.462 0.469 0.458 0.469 0.467 0.473
hard prompt 3 5 10 15
MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.378 0.405 0.376 0.402 0.376 0.402 0.375 0.400
munichCarbon 0.236 0.350 0.239 0.352 0.242 0.354 0.236 0.350
chinaCarbon 0.464 0.474 0.462 0.469 0.466 0.469 0.460 0.467
d_model 32 64 128 256
MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.376 0.402 0.378 0.402 0.378 0.402 0.381 0.403
munichCarbon 0.239 0.352 0.236 0.349 0.241 0.352 0.238 0.350
chinaCarbon 0.462 0.469 0.467 0.473 0.470 0.477 0.455 0.468
d_ff 32 64 128 256
MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.381 0.405 0.380 0.406 0.376 0.402 0.380 0.400
munichCarbon 0.239 0.352 0.243 0.354 0.239 0.352 0.239 0.351
chinaCarbon 0.482 0.483 0.460 0.469 0.462 0.469 0.464 0.471
Ablation Experiment
As shown in Table 4, we conduct an ablation study on multiple datasets (ETTh1, ETTm1, munichCarbon, chinaCarbon, worldCarbon) to analyze the effects of soft prompts (SP), hard prompts (HP), cross-modal alignment (CMA), and LoRA. All experiments follow the long-term forecasting setting, using four different prediction horizons, and the reported results are averaged across these settings. The results indicate that all these modules activate the LLM’s time series forecasting capabilities to varying degrees. Compared to the prompting engineering, LoRA and cross-modal alignment have a more significant impact on the LLM’s forecasting ability. Additionally, we replace GPT-2 with the Qwen-3 0.6B model, which lead to a significant performance drop. We hypothesize that this is because larger-parameter models struggle to effectively handle continuous-valued time-series data. This observation aligns with findings in the time-series foundation model literature: most published time-series foundation models are relatively small, typically in the few-hundred-million parameter range or even smaller (Team 2025). This also indicates that in future research, we should focus more on the design of model architecture and the tuning of model parameters, rather than solely on prompt engineering. Furthermore, greater attention should be paid to foundation models and small-scale LLMs (Liu et al. 2025b; Shi et al. 2025).
Modal Analysis
We experiment with the settings for soft prompt length, hard prompt length, $d_{model}$, and $d_{ff}$ on multiple datasets (ETTh1, munichCarbon, chinaCarbon). The hard prompt length refers to the number of tokens retained after the hard prompt is encoded by the LLM. Due to the relatively large number of tokens, we truncate it to mitigate the curse of dimensionality. This approach is supported by existing literature, which indicates that the key information is predominantly located in the latter part (Liu et al. 2025a). Here, $d_{model}$ represents the intermediate dimension of the patch reprogram module, while $d_{ff}$ is the intermediate dimension of the projection module.
As shown in the Table 5, the soft prompt length should neither be too short nor too long. Since soft prompts are designed to guide the LLM’s behavior, an excessively long sequence of tokens is unnecessary. In contrast, hard prompts are constructed by textualizing time-series data, so retaining more tokens is generally better, as truncation inevitably leads to information loss. However, considering the trade-off between efficiency and performance, we opt to keep only the most informative segment. For $d_{ff}$ and $d_{model}$, we employ a grid search to select the optimal parameter combination. Ultimately, for all datasets, we default to $d_{ff}=128$ and $d_{model}=32$.
Conclusion
This paper proposes Time-Prompt to unlock the potential of LLMs in time series forecasting. This method employs a dual-path prompting mechanism. It utilizes soft prompts to guide the reasoning patterns of the LLM and enhances representations with hard prompts and time series reprogramming. Subsequently, it bridges the modality gap between time series and text through semantic space embedding and cross-modal alignment. Finally, the LLM’s capability for processing continuous values is enhanced through LoRA fine-tuning. Extensive experiments on multiple real-world datasets validate the effectiveness of Time-Prompt in time series forecasting.
Despite preliminary advancements, we find that the time series forecasting capabilities of LLMs remain inadequate. Whether the discrete reasoning abilities of language models can be effective for time series forecasting is still an open question. We argue that this fundamental limitation stems from the text-centric training paradigm of LLMs, which inherently struggles to adapt to forecasting tasks. Moreover, according to the experiments in this paper, we find that small-scale LLMs are more effective than LLMs, and fine-tuning is more effective than prompt engineering in activating predictive capabilities—especially in few-shot scenarios. So our future work will be dedicated to developing time series foundation models or exploring the time-series forecasting capabilities of small-scale LLMs.